40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14815     Entry time: Mon Jul 29 13:32:56 2019
Author: gautam 
Type: Update 
Category: Loss Measurement 
Subject: Loss measurement PD installed in AS path 

[yehonathan, gautam]

  • we placed a PDA520 photodiode in the AS beampath, so AS110 and AS55 no longer see any light.
  • ITMX and ETMX were misaligned (since the plan is to measure the Y arm loss).
  • The PDA520 and MC2 transmission are currently going to the Y arm ALS beat channels in the DAQ system. Unfortunately, we have no control over the whitening gains for these channels because of the c1iscaux2 situation.
ELOG V3.1.3-